IF angle KNL = 10y-12, and angle JNM = 8y+4
Then
Find Y.
Find the measure of angle JNK.
.__ K ____________________J.
____. _________________.
_______. ___________.
_________.___N____.
____________. .__
_________.________.
_______._______________.
____L .___________________.M

Answer by Edwin McCravy(20081)



Angles KNL and JNM are vertical angles so they are equal so

     10y - 12 = 8y + 4
           2y = 16
            y = 8

So angle KNL = 10y - 12 = 10(8) - 12 = 80 - 12 = 68°

and, as a check

Angle JNM = 8y + 4 = 8(8) + 4 = 64 + 4 = 68°

-----------------

So now we have  



and we want to find the angle JNK indicated by the green arc.

Angle KNL and angle JNK are supplementary, so 

Angle JNK  = 180° - Angle KNL

So

Angle JNK = 180° - 68° = 112°



Edwin